What is the percentage of 320 out of 359?
Answer by stanbon(75887)

What is the percentage of 320 out of 359?
---
x/100 = 320/359
---
x = 100(320/359)
---
x = 100*0.8914
---
x = 89.14%
============
Cheers,
Stan H.